OpenStudy (anonymous):

a bookcase is to be constructed with three shelves as shown.the height of the bookcase is 4 feet longer than the length of a shelf and only 20ft of lumber is available.what should be the width and height of the bookcase.

OpenStudy (anonymous):

how can i show you the picture

OpenStudy (anonymous):

how many total horizontal pieces are there? This matters.

OpenStudy (anonymous):

In othere words does is there just three shelves or 3 shelves plus top or three shelves plus top and bottom.

OpenStudy (anonymous):

4 if you include the top

OpenStudy (anonymous):

length of shelf = x Needed total lengths of shelf = 3x (three shelves) Needed total lengths (with top and bottom of bookcase) =4x Height of bookcase = x + 4 Needed total length for height (right and left of bookcase) = 2x +8 Total amount of wood needed = 6x +8 = 20 ft (amount you have) Width of bookcase = x = 2 ft Length of bookcase = x +4 = 6 ft

OpenStudy (anonymous):

x = length of shelf x+4 = height 2(x+4)+ 4x = 20 x=shelf width = 2 feet height = 6 feet